The sum of three consecutive integers times four is 180.(I know the integers are 14 15 16 but how do you get that? I need equation and solution please.)!

Answer by checkley77(12569)

4(x+x+1+x+2)=180
4(3x+3)=180
12x+12=180
12x=180-12
12x=168
x=168/12
x=14 answer for the first integer.
14=1=15 for the second one
14=2=16 for the last one
proof
4(14+15+16)=180
4*45=180
180=180